A rectangle is drawn so that the width is 5 feet shorter than the length. The area of the rectangle is 36 square feet. Find the length of the rectangle.
Answer by Cromlix(4381)

Hi there,
Make length = x
width = x - 5
Area = length x width
36 = x(x - 5)
36 = x^2 - 5x
x^2 - 5x - 36 = 0
(x - 9)(x + 4) = 0
x + 4 = 0
x = -4 disregard as -ve
x - 9 = 0
x = 9
Length = 9 feet
Width 4 feet
Hope this helps :-)